OpenStudy (anonymous):

Find the area between a curve. X=0 ,x=siny, y=pi/4, y=3pi/4 Anyways I set it up and got -cosy as my antiderivative. And when I plug in the boundaries I get sqrt2/2 both times to equal zero. My book says the answer is sqrt2 what am I doing wronf?

OpenStudy (irishboy123):

you should get \(\large [-cos y]_{\pi/4}^{3\pi/4}\)

OpenStudy (irishboy123):

so i assume algebraic error (minus signs) as you seem to have got that..... \(-(-\frac{1}{\sqrt{2 } } )-(-\frac{1}{\sqrt{2 } })\)

OpenStudy (amistre64):

-cos(3pi/4) +cos(pi/4) --1/sqrt2 + 1/sqrt2 = 2/sqrt2 = sqrt2 another way to see it is: 1/sqrt2 = sqrt2/2

OpenStudy (amistre64):

yeah, he is missing the minus sign to me as well